a and b are equal partners with capitals 2,00,000 and 1,00,000 respectively. as per deed, they are allowed a interest @8% on capital . during the year the firm earned a profit of rs 12,000. interest in capital allowed to a will be

1
Expert's answer
2021-11-22T10:08:43-0500

Solution:

A = Interest on capital = "8\% \times 200,000 = 16,000"

B = Interest on capital = "8\% \times 100,000 = 8,000"

Total interest on capital = 16,000 + 8,000 = 24,000

 

Profit before interest = 12,000

 

Interest on capital allowed:

A = "\frac{12,000\times16,000}{24,000} = 8,000"

Interest in capital allowed for A = 8,000